OpenStudy (anonymous):

Write an equation for a line in slope-intercept form that passes through the points (3, 4) and (1, -2).

OpenStudy (anonymous):

m=(y1-y2)/(x1-x2) m=(4--2)/(3-1)=6/2=3 Thus y=3x+b (1,-2) -2=3+b ->b=-5 So, y=3x-5

OpenStudy (anonymous):

Thank you!